OpenStudy (dj3strella):

Question 1.1. Select the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2 y – 9 = 2(x – 5) y – 5 = 2(x + 9) y + 9 = 2(x – 5) y – 9 = 2(x + 5)

OpenStudy (anonymous):

y- y1 = m(x-x1) m = slope

OpenStudy (anonymous):

(x1, y1) (5, 9) plug the points